A sports car accelerates in third gear from 48.3 km/h (about 30 mi/h) to 80.5 km/h (about 50 mi/h) in 3.70 s. a) What is the average acceleration of this car in mi/s? b) If the car maintained this acceleration, how fast would it be moving one second later?
1
Expert's answer
2021-10-26T10:51:48-0400

a) Average acceleration

"A_{av}=\frac{dv}{dt}"


"=\frac{80.5-48.3}{3.6×3.7} =2.42m/s^2"

3.6 is change from km/h to m/s


b)

"V=Vi+a\u00d7t"


"=\frac{80.5}{3.6}+2.42×1.0=24.781m/s=89.211km/h"